OpenStudy (jerudavis):

WILL MEDAL Pia can make 1/2 a party favor in 1/4 h. _______ How many party favors can she make in 1 h?

OpenStudy (jerudavis):

@mathmale

OpenStudy (jerudavis):

@Kalayn321

OpenStudy (anonymous):

2

OpenStudy (mathmale):

You are given a rate of change (how many party favors can be made in a period of time). This rate is (1/2) party favor in (1/4) hour, or \[\frac{ \frac{ 1 }{ 2 }favor }{ \frac{ 1 }{ 4 }hr }\]

OpenStudy (mathmale):

First, please simplify this. How would you divide 1/4 into 1/2? Please show your work.

OpenStudy (jerudavis):

so @Kalayn321 is not right @mathmale

OpenStudy (mathmale):

I'd much prefer to focus on this math instead of whether another OpenStudy member is right or not.

OpenStudy (mathmale):

Want to see YOUR work.

OpenStudy (jerudavis):

im asking because I have to hurry

OpenStudy (mathmale):

I sympathize, but I do not work that way.

OpenStudy (mathmale):

Show how you'd divide 1/4 into 1/2, please.

OpenStudy (jerudavis):

Kalayn are u sure

OpenStudy (briellshaw):

SO yawl met???

OpenStudy (anonymous):

If she makes 1/2 in 1/4 of an hour which wouuld be 15. How many would she make in 1 hour.

OpenStudy (anonymous):

yes im sure

OpenStudy (mathmale):

What are the units of measurement? \[\frac{ \frac{ 1 }{ 2 }favor }{ \frac{ 1 }{ 4 }hr }=?\]

OpenStudy (anonymous):

Did you get it right?

OpenStudy (jerudavis):

yep

OpenStudy (anonymous):

this guy ~__~

OpenStudy (mathmale):

So, where's your final answer? Can you defend this answer (explain why you believe it's correct)?